A mole weighs about 1.5 ounces. To dig up our yards it has 10 claws on its front feet. It actually has two eyes but they can only sense light and dark - it really can't see.

1. If you had three moles of moles what would be their mass in kilograms?

2. How many total claws would there be in the three moles of moles?

3. How many total eyes in five moles of moles?
Answer #1

A mole(insects) weights = 1.5 ounces = 0.04252 kg

1 mole = 6.022 * 10 23 items

3 moles of moles = 3 * 6.022*10 23 moles

Wt of 3 moles of moles = 0.04252 * 3 * 6.022*10 23 kg

= 7.6816 * 1022 kg

2. No of claws per mole(insect) = 10

No of claws 3 moles of moles = 10 * 3 * 6.022 * 1023

= 1.8066 * 1025 claws

3. No. of eyes per mole(insect) = 2

No. Of eyes in 5 moles of moles = 2 *5* 6.022*1023

= 6.022*10 24 eyes